OpenStudy (anonymous):

sin^-1(sin(17pi/19))

hartnn (hartnn):

\(\huge sin^{-1}(sin \:a)=a\)

hartnn (hartnn):

so what will be sin^-1(sin(17pi/19)) =?

OpenStudy (anonymous):

The answer is \[\frac{ 2\pi }{19}\] But I don't understand how or why.

OpenStudy (anonymous):

is it because it's not within \[\frac{ -\pi }{ 2 } and \frac{ \pi }{ 2 }\] ?

hartnn (hartnn):

yes, 17pi/19 is not in that range, so u need to add or subtract 2\(\pi\) to bring the angle in that range.

OpenStudy (anonymous):

Did you mean just pi?

OpenStudy (watchmath):

the arcsin is defined on the interval \((-\pi/2,\pi/2)\) so we need to do something first with this problem. \(\sin^{_1}(\sin 17\pi/19)=\sin^{-1}(\sin 2\pi/19)=2\pi/19\)

OpenStudy (anonymous):

since it's in QII we use theta minus pi right?

OpenStudy (watchmath):

\(\sin \theta=\sin (\pi-\theta)\)

hartnn (hartnn):

u bring it in 1st quadrant by using that formula ^ so sin(17pi/19)= sin(pi-17pi/19)

OpenStudy (anonymous):

oops, meant pi minus theta. Thank you both!